the boiling point of an aqueous solution of a non volatile solute is  100.150C .what is the freezing point of an aqueous solution obtained by diluting the above solution with an equal volume of water? the values of Kb and Kf for water are 0.5120C and 1.860C K molal -1

Grade:12

3 Answers

Vikas TU
7 years ago
For boiling point,
Tb = i * kb * M
100.15 + 273 = 1*0.512*M
from here calculate M.
After adding 500 ml water,
molarity becomes,
M x V = M’ x 2V
M’ = M/2
Then apply,
Tf = 1*kf*M’
ankit singh
3 years ago
this the attcahment of the question that u want to know
ankit singh
manit bhopal
thope the answer is useful to u keep it up work hard to fullfill your drea
ankit singh
3 years ago
 
 
For boiling point,
Tb = i * kb * M
100.15 + 273 = 1*0.512*M
from here calculate M.
After adding 500 ml water,
molarity becomes,
M x V = M’ x 2V
M’ = M/2
Then apply,
Tf = 1*kf*M’